### v1

@@ 9 Discussion
In this paper we have introduced the concept Spherical Beltrami Differential, which has one to one correspondence with the equivalence class of spherical diffeomorphism up to the group of conformal mappings. Following this, we used Spherical Beltrami Network (SBN) to establish a neural optimization framework BOOST for spherical parameterization problems, which achieves optimization of spherical Beltrami differentials in a two-chart manner. Across multiple distinct testbeds including cortical surface alignment (Sec.8.1), extreme distortions (Sec.8.2), and hybrid intensity‐landmark matching (Sec.8.3), BOOST matched or beat both classical numerical algorithms. Here, we try to analyze several potential factors that underlie this excellence:
@@
    Power of SBN BOOST’s effectiveness stems directly from the fidelity of LSQC’s neural surrogate, the Spectral Beltrami Network (SBN). By closely approximating the LSQC objective, SBN ensures that for any admissible Beltrami coefficient (BC) field, the predicted mapping tracks the true LSQC solution and usually remains diffeomorphic, thereby realizing the deformation prescribed by the input BC. This high-quality differentiable approximation gives BOOST a smooth, well-conditioned optimization landscape in Beltrami space, enabling gradient-based updates and suited for general spherical parameterization problems even some cases that explicit gradient on the deformation mapping is hard to compute.
@@
    Free-boundary optimization Traditional numerical algorithms usually impose hard boundary constraints and require mappings after each iteration to be non-overlapping, which shrinks the admissible solution space and increases the risk of getting trapped in sub-optimal local minima. SBN inherits the free-boundary perspective of LSQC energy, which provides a broader search space of mappings. Additionally, BOOST permits temporary foldings and boundary drift during optimization, which allows the optimizer bypass poor minima and folding would be removed when the folding loss dominates.
@@
    Direct control on conformality While classical methods often minimize angle distortion indirectly(e.g. Laplaician smoothing) and deep learning approaches can control angle distortion only through penalizing the output mapping, our BOOST, by optimizing Beltrami coefficients directly via smoothness and magnitude penalties, achieves fine‐grained angular‐distortion control.
The excellence of SBN is not just that it “uses deep learning,” but that it faithfully encodes a mathematically well-behaved energy whose properties—existence, uniqueness, invariance, and resolution-independence, which translate directly into superior performance. By these, SBN achieves what neither classical solvers nor generic neural networks can: produce accurate surface mappings with well-controlled distortion and generalize across resolutions and different applications.
Though SBN represents a substantial advance and BOOST outperforms in applications, several pragmatic gaps remain:
@@
    Not totally scale-free. Currently, to handle ultra-high-resolution meshes, we interpolate the solution from a lower-resolution “standard” grid, which is supported by the resolution-independence property of LSQC. However, this property is built on the assumption that beltrami coefficients on the three sub-triangles equal to that of the triangle before split. This condition might be inappropriate in high-curvature or feature-dense regions, and can introduce interpolation artifacts not faithfully reflecting fine-mesh geometry. An ideal remedy is a truly resolution-invariant architecture (e.g. continuous operators or adaptive patch hierarchies) that natively ingests Beltrami data and outputs deformation at arbitrary granularity, and that would be our future working direction.
@@
    Limited to genus-0 topology. Our current framework BOOST assumes genus-0 (spherical) topology. Extending to higher-genus or open surfaces is straightforward in principle—by redefining shape of parametrization domain as well as chart overlaps, and this is also left to future work.
@@ 10 Conclusions
In this work, we introduced a novel concept Spherical Beltrami Differential and build correspondence between Spherical Beltrami Differentials and spherical self-quasiconformal mappings. Following this, we proposed a two-chart optimization framework BOOST tailored to complex genus-0 surface mapping problems. As a key factor in BOOST, SBN can produce free-boundary quasiconformal mappings accurately following the input BCs. Besides, through designing three auxilliary losses $\mathcal{L}_{\text{bm}},\mathcal{L}_{\text{bs}}\text{ and }\mathcal{L}_{\text{folding}}$, BOOST resolves gluing issues between two hemispheres, and achieving smooth transitions across hemispherical boundaries. These capabilities establish BOOST as a robust and versatile framework for spherical parameterization problems. The excellence of BOOST has been demonstrated through extensive experiments and applications across three key categories: Landmark matching ( Sec.8.1, 8.2), Intensity matching (e.g., Sec.8.1) and Hybrid matching (e.g., Sec.8.3).
In conclusion, it is the first time that we achieve optimization over surface Beltrami differentials and the outstanding performance of BOOST implies the potential of exploring optimization over surface Beltrami Differentials on general surfaces.

### v2

@@ 8 Discussion
In this paper we have introduced the concept of the Spherical Beltrami Differential, which has a one-to-one correspondence with the equivalence class of spherical homeomorphisms up to the group of conformal mappings. Following this, we used the Spherical Beltrami Network (SBN) to establish a neural optimization framework BOOST for spherical parameterization problems, which achieves optimization of spherical Beltrami differentials in a two-chart manner. Across multiple distinct testbeds including cortical surface alignment, extreme distortions, and hybrid intensity‐landmark matching, BOOST matched or beat both classical numerical algorithms. Here, we try to analyze several potential factors that underlie this excellence:
@@
    Power of SBN. The effectiveness of BOOST stems directly from the capacity of the LSQC neural surrogate, the Spectral Beltrami Network (SBN). By closely approximating the LSQC objective, SBN ensures that for an admissible Beltrami coefficient (BC) field, the predicted mapping tracks the true LSQC solution and usually remains bijective, thereby realizing the deformation prescribed by the input BC. This high-quality differentiable approximation gives BOOST a smooth, well-conditioned optimization landscape in the Beltrami space, enabling gradient-based updates and suited for general spherical parameterization problems even some cases that explicit gradient on the deformation mapping is hard to compute.
@@
    Free-boundary optimization. Traditional numerical algorithms usually impose hard boundary constraints, which shrinks the admissible solution space and increases the risk of getting trapped in sub-optimal local minima. SBN inherits the free-boundary perspective of LSQC energy, which provides a broader search space of mappings. Moreover, the discrete choice of pinpoint required in LSQC becomes differentiable in SBN, i.e. BOOST automatically learns what is the most appropriate pinpoint condition, and thus BOOST can be applied in landmark-free mapping problems.
@@
    Direct control on conformality. While classical methods often minimize angle distortion indirectly (e.g. Laplacian smoothing) and deep learning approaches can control angle distortion only through penalizing the output mapping, our BOOST, by regularizing the Beltrami differential directly via smoothness and magnitude penalties, achieves fine‐grained angular‐distortion control.
The excellence of SBN lies not just in its use of deep learning, but that it faithfully encodes a mathematically well-behaved energy whose properties—existence, uniqueness, invariance, and resolution-independence, which translate directly into superior performance. By these, SBN achieves what neither classical solvers nor generic neural networks can: produce accurate surface mappings with well-controlled distortion and generalize across resolutions and different applications. Though SBN represents a substantial advance and BOOST outperforms in applications, several pragmatic gaps remain:
@@
    Not totally mesh-free. Currently, to handle ultra-high-resolution meshes, we interpolate the solution from a lower-resolution “standard” grid, which is supported by the resolution-independence property of LSQC. However, this property is built on the assumption that Beltrami coefficients on the three sub-triangles equal to that of the parent triangle before splitting. This condition might be inappropriate in high-curvature or feature-dense regions, and can introduce interpolation artifacts that do not faithfully reflecting fine-mesh geometry. An ideal remedy is a truly resolution-invariant architecture (e.g. continuous operators or adaptive patch hierarchies) that naturally receives Beltrami data and outputs deformation at arbitrary granularity, and that would be our future working direction.
@@
    Limited to genus-0 topology. Our current framework BOOST assumes genus-0 (spherical) topology. Extending to higher-genus or open surfaces is straightforward in principle—by redefining shape of parameterization domain as well as chart overlaps, and this is also left to future work.
@@ 9 Conclusions
In this work, we introduced the novel concept of the Spherical Beltrami Differential and built a correspondence between Spherical Beltrami Differentials and spherical self-quasiconformal mappings. Following this, we proposed a two-chart optimization framework BOOST tailored to complex genus-0 surface mapping problems. As a key factor in BOOST, SBN can produce free-boundary quasiconformal mappings accurately following the input BCs. By designing three auxiliary losses $\mathcal{L}_{\text{bm}},\mathcal{L}_{\text{bs}}\text{ and }\mathcal{L}_{\text{folding}}$, BOOST resolves gluing issues between two hemispheres, and achieves smooth transitions across hemispherical boundaries. These capabilities establish BOOST as a robust and versatile framework for spherical parameterization problems. The excellence of BOOST has been demonstrated through extensive experiments and applications across three key categories: Landmark matching, Intensity matching and Hybrid matching registration.
To the best of our knowledge, this is the first work to achieve optimization over surface Beltrami differentials, and the outstanding performance of BOOST implies the potential of exploring optimization over surface Beltrami Differentials on general surfaces.